how do i solve
8x-(3x+4)=6
the solution is x=

Answer by Nieleya(22)

8x-(3x+4)=6
the solution is x=
8x - 3x - 4 = 6
5x - 4 = 6
5x = 6 + 4
5x = 10
x = 10/5
x = 2